OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4.

OpenStudy (anonymous):

a) 2 b) 3 c) All real numbers d) No solution

OpenStudy (aum):

\[ -2(x + 3) = -2(x + 1) - 4 \\ -2x - 6 = -2x - 2 - 4 \\ -2x - 6 = -2x - 6 \]The LHS and RHS are identical no matter what the value of x is. Therefore, the solution is: All real numbers.

OpenStudy (anonymous):

thankssss